### v3

# Shivaram Gopal    SM Ferdous    Hemanta Maji    Alex Pothen

Abstract
We describe a parallel approximation algorithm for maximizing monotone submodular functions subject to hereditary constraints on distributed memory multiprocessors. Our work is motivated by the need to solve submodular optimization problems on massive data sets, for practical contexts such as data summarization, machine learning, and graph sparsification.
Our work builds on the randomized distributed RandGreedi algorithm, proposed by Barbosa, Ene, Nguyen, and Ward (2015). This algorithm computes a distributed solution by randomly partitioning the data among all the processors and then employing a single accumulation step in which all processors send their partial solutions to one processor. However, for large problems, the accumulation step exceeds the memory available on a processor, and the processor which performs the accumulation becomes a computational bottleneck.
Hence we propose a generalization of the RandGreedi algorithm that employs multiple accumulation steps to reduce the memory required. We analyze the approximation ratio and the time complexity of the algorithm (in the BSP model). We evaluate the new GreedyML algorithm on three classes of problems, and report results from large-scale data sets with millions of elements. The results show that the GreedyML algorithm can solve problems where the sequential Greedy and distributed RandGreedi algorithms fail due to memory constraints. For certain computationally intensive problems, the GreedyML algorithm is faster than the RandGreedi algorithm. The observed approximation quality of the solutions computed by the GreedyML algorithm closely matches those obtained by the RandGreedi  algorithm on these problems.
keywords:
Combinatorial optimization, submodular functions, distributed algorithms, approximation algorithms, data summarization.
category:
[…]
1 Introduction
We describe GreedyML, a parallel approximation algorithm for maximizing monotone submodular functions subject to hereditary constraints on distributed memory multiprocessors. GreedyML is built on an earlier distributed approximation algorithm, which has limited parallelism and higher memory requirements. Maximizing a submodular function under constraints is NP-hard, but a natural iterative Greedy algorithm exists that selects elements based on the marginal gain (defined later) and is $(1-1/\mathrm{e})\approx 0.63$-approximate for cardinality constraints and $1/2$-approximate for matroid constraints; here $\mathrm{e}$ is Euler’s number.
Maximizing a submodular function (rather than a linear objective function) promotes diversity in the computed solution since at each step the algorithm encourages an element with the least properties in common with the current solution set. A broad collection of practical problems are modeled using submodular functions, including data and document summarization [21], load balancing parallel computations in quantum chemistry [7], sensor placement [5], resource allocation [26], active learning [9], interpretability of neural networks [6], influence maximization in social networks [11], diverse recommendation [4]. Surveys discussing submodular optimization formulations, algorithms, and computational experiments include Tohidi et al. [27] and Krause and Golovin [12].
Our algorithm builds on the RandGreedi  framework [2], a state-of-the-art randomized distributed algorithm for monotone submodular function maximization under hereditary constraints, which has an approximation ratio half that of the Greedy algorithm. The RandGreedi algorithm randomly partitions the data among all the processors, runs the standard Greedy algorithm on each partition independently in parallel, and then executes a single accumulation step in which all processors send their partial solutions to one processor. However, this accumulation step could exceed the memory available on a processor when the memory is small relative to the size of the data, or when solutions are large. Additionally, the accumulation serializes both the computation and communication and is a bottleneck when scaled to many machines.
Our GreedyML algorithm brings additional parallelism to this step and can lower the memory and running time by introducing hierarchical accumulation organized through an accumulation tree. Similar to RandGreedi, we randomly partition the data among all the processors, which constitute the leaves of the accumulation tree. We merge partial solutions at multiple levels in the tree, and the final solution is computed at the root. We prove that the GreedyML algorithm has a worst-case expected approximation guarantee of $(\alpha\>b)/(m+b)$, where $\alpha$ is the approximation guarantee for the Greedy algorithm, $b$ is the branching factor, and $m$ is the number of leaves in the accumulation tree. Using the BSP model, we also analyze the time and communication complexity of the GreedyML  and RandGreedi algorithms and show that the former has lower computation and communication costs than the latter.
We evaluate the parallel algorithms on three representative and practical submodular function maximization problems: maximum $k$-set cover, maximum $k$-vertex dominating set in graphs, and exemplar-based clustering (modeled by the $k$-medoid problem). We experiment on large data sets with millions of elements that exceed the memory constraints (a few GBs) on a single processor, and demonstrate how to choose the accumulation tree to have more levels to adapt to the small memory available on a processor. This strategy also enables us to solve for larger values of the parameter $k$ in the problems discussed above, which corresponds to the size of the solution sought. We also show that the number of function evaluations on the critical path of the accumulation tree, and hence the run time, could be reduced when the parallel algorithm is employed. In most cases, we find the quality of the computed solutions by our GreedyML closely matches those obtained by the RandGreedi  algorithm on these problems despite having a worse expected approximation guarantee.
2 Background and Related Work
Submodular functions
A set function $f\colon 2^{W}\to\mathbb{R}^{+}$ defined on the power set of a ground set $W$ is submodular if it satisfies the diminishing marginal gain property. That is,
$f(X\cup\{w\})-f(X)\geqslant f(Y\cup\{w\})-f(Y),\text{ for all }X\subseteq Y\subseteq W\text{ and }w\in W\setminus Y.$
A submodular function $f$ is monotone if for every $X\subseteq Y\subseteq W$, we have $f(X)\leqslant f(Y)$. The constrained submodular maximization problem is defined as follows.
$\max f(S)\text{ subject to }S\in\mathcal{C},\text{where }\mathcal{C}\subseteq 2^{W}\text{ is the family of feasible solutions.}$.
We consider hereditary constraints: i.e., for every set $S\in\mathcal{C}$, every subset of $S$ is also in $\mathcal{C}$. The hereditary family of constraints includes various common ones such as cardinality constraints ($\mathcal{C}=\{A\subseteq W:|A|\leqslant k\}$) and matroid constraints ($\mathcal{C}$ corresponds to the collection of independent sets of a matroid).
Lovász extension
For the analysis of our algorithm, we use the Lovász extension [19], a relaxation of submodular functions. A submodular function $f$ can be viewed as a function defined over the vertices of the unit hypercube, $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{+}$, by identifying sets $V\subseteq W$ with binary vectors of length $w=|W|$ in which the $i^{\text{th}}$ component is $1$ if $i\in V$, and 00 otherwise. The Lovász extension [19] $\widehat{f}:[0,1]^{w}\rightarrow\mathbb{R}^{+}$ is a convex extension that extends $f$ over the entire hypercube and given by, $\widehat{f}(x)=\mathop{\mathbb{E}}\limits_{\theta\in\mathcal{U}[0,1]}\left[\;f\left(\{i:x_{i}\geqslant\theta\}\right)\;\right].$ Here, $\theta$ is uniformly random in $[0,1]$. The Lovász extension $\widehat{f}$ satisfies the following properties [19]:
  1. 1.
    $\widehat{f}(1_{S})=f(S)$, for all $S\subseteq V$ where $1_{S}\in[0,1]^{w}$ is a vector containing 1 for the elements in $S$ and 00 otherwise,
  2. 2.
    $\widehat{f}(x)\ \text{is convex}$, and
  3. 3.
    $\widehat{f}(c\cdot x)\geqslant c\cdot\widehat{f}(x),\text{for any}\ c\in[0,1]$.
An $\alpha$-approximation algorithm ($\alpha\in[0,1)$) for constrained submodular maximization produces a feasible solution $S\subseteq W$, satisfying $f(S)\geqslant\alpha\cdot f(S^{*})$, where $S^{*}$ is an optimal solution.
2.1 Related Work
Greedi and RandGreedi.
The iterative Greedy algorithm for maximizing constrained submodular functions starts with an empty solution. Given any current solution $S$, an element is feasible if it can be added to the solution without violating the constraints. In each iteration, the Greedy algorithm on a dataset $V$ chooses a feasible element $e\in V$ that maximizes the marginal gain, $f(S\cup\{e\})-f(S)$, w.r.t. the current solution $S$. The algorithm terminates when the maximum marginal gain is zero or all feasible elements have been considered.
We now discuss the Greedi and RandGreedi algorithms, which are the state-of-the-art distributed algorithms for constrained submodular maximization. The Greedi algorithm [21] partitions the data arbitrarily on available machines, and on each machine, it runs the Greedy algorithm independently to compute a local solution. These solutions are then accumulated to a single global machine. The Greedy algorithm is again executed on the accumulated data to get a global solution. The final solution is the best solution among all the local and global solutions. For a cardinality constraint, where $k$ is the solution size, the Greedi algorithm has a worst-case approximation guarantee of $1/\Theta(\min(\sqrt{k},m))$, where $m$ is the number of machines.
Although Greedi performs well in practice [21], its approximation ratio is not a constant but depends on $k$ and $m$. Improving on this work, Barbosa et al. [2] proposed the RandGreedi algorithm, which partitions the data uniformly at random on machines and achieves an expected approximation guarantee of $\frac{1}{2}(1-1/\mathrm{e})$ for cardinality and $1/4$ for matroid constraints. In general, it has an approximation ratio of $\alpha/2$ where $\alpha$ is the approximation ratio of the Greedy algorithm used at the local and global machines. We present the pseudocode of RandGreedi framework in Algorithm 1 in Appendix B. Note that for a cardinality constraint, both Greedi and RandGreedi perform $O(nk(k+m))$ calls to the objective function and communicate $O(mk)$ elements to the global machine where $n$ is the number of elements in the ground set, $m$ is the number of machines, and $k$ is solution size.
Both Greedi and RandGreedi require a single global accumulation from the solutions generated in local machines that can quickly become dominating since the runtime, memory, and complexity of this global aggregation grows linearly with the number of machines. We propose to alleviate this by introducing a hierarchical aggregation strategy that maintains an accumulation tree. Our GreedyML framework generalizes the RandGreedi from a single accumulation to a multi-level accumulation. The number of partial solutions to be aggregated depends on the branching factor of the tree, which can be a constant. Thus, the number of accumulation levels grows logarithmically with the number of machines, and the total aggregation is not likely to become a memory, runtime, and communication bottleneck with the increase in the number of machines. We refer to Appendix C for the detailed complexity comparisons of the RandGreedi and our GreedyML algorithm.
Other work. Early approaches on distributed submodular maximization includes the $1/(2+\varepsilon)$-approximate Sample and Prune algorithm by Kumar et al. [15], which requires $O(1/\delta)$ rounds assuming $O(kn^{\delta}\>\log{n})$ memory per machines. Here, $\delta>0$ is a user parameter. Greedi [21] and RandGreedi [2] are shown to be more efficient in practice than the Sample and Prune algorithm.
More recent distributed approaches [3, 22, 23] use the multi-linear extension to map the problem into a continuous function. They typically perform a gradient ascent on each local machine and build a consensus solution in each round, which improves the approximation factor to $(1-1/\mathrm{e})$. However, we do not believe that these approaches are practical since they involve expensive gradient computations (could be exponential-time). Most of these algorithms are not implemented, and the one reported implementation solves problems with only a few hundred elements in the data set [23].
3 Description of Our Algorithm
We describe and analyze our algorithm that generalizes the RandGreedi algorithm from a single accumulation step to multiple accumulation steps. Each accumulation step corresponds to a level in an accumulation tree, which we describe next. We assume that there are $m$ machines identified by the set of ids: $\{0,1,\ldots,m-1\}$.
Accumulation tree. An accumulation tree ($T$) is defined by the number of machines ($m$), and branching factor ($b$). It has the same structure as a complete $b$-ary tree with $m$ leaves, which means all the leaves are at the same depth. The tree nodes correspond to processors and the corresponding subset of data accessible to them. The edges of the tree determine the accumulation pattern of the intermediate solutions. The final solution is generated on the root node of $T$. Thus, the branching factor $b$ of the tree indicates the maximum number of nodes that transmit data to its parent. The number of accumulation levels (i.e., one minus the height of the tree), denoted by $L$, is $\lceil\log_{b}{m}\rceil$.
To uniquely identify a node in the tree, we assign an identifier $(\ell,id)$ to each node of $T$, where $\ell$ represents the accumulation level of the node and $id$ represents the machine identifier corresponding to the node. The $id$ for each leaf node is the identifier of the machine to which the leaf node corresponds. All the leaf nodes are at the zeroth level. Each internal node receives the lowest $id$ of its children, i.e., any node $(\ell,i)$ has node $(\ell+1,\lfloor i/b^{\ell+1}\rfloor*b^{\ell+1})$ as the parent. Therefore, the root node will always reside at level $L$ with $id=0$. Also, we characterize an accumulation tree $T$ by the triple $T(m,L,b)$, where $m$ is the number of leaves (machines), $L$ is the number of levels, and $b$ is the branching factor.
{forest}
[…]
Figure 1: An accumulation tree with $L=2$ levels, $m=b^{2}$ machines, and a branching factor $b$. Each node has a label of the form $(\ell,id)$. Here there are $b$ nodes as children at each level, but when there are fewer than $b^{L}$ leaf nodes, then the number of children at levels closer to the root may be fewer than $b$.
Figure 1 shows an example of a generic accumulation tree with $b^{2}$ leaves and branching factor $b$. The number of accumulation levels is the level of the root. Here we have $L=\lceil\log_{b}b^{2}\rceil=2$. We also show a few realizations of different accumulation trees in Figure 6 in the Appendix A.
Observe that the $id$ parameter remains the same in multiple nodes that are involved in computations at multiple levels. For our analysis, we keep the branching factor constant across all levels.
Randomness. The randomness in the algorithm is only in the initial placement of the data on the machines, and we use a random tape to encapsulate this. The random tape $r_{W}$ has a randomized entry for each element in $W$ to indicate the machine containing that element. Any expectation results proved henceforth are over the choice of this random tape. Moreover, if the data accessible to a node is $V$, we consider the randomness over just $r_{V}$. Whenever the expectation is over $r_{V}$, we denote the expectation as $\mathbb{E}_{V}$.
$\mbox{\sc GreedyML}(\ell,id)=$
$\begin{cases}\textsc{Greedy}(P_{id})&\ell=0\\ \arg\max\begin{cases}\textsc{Greedy}\left(\bigcup\limits_{i\in\{0,1,\dotsc,b-1\}}\mbox{\sc GreedyML}\left(\ell-1,id+i\cdot b^{\ell-1}\right)\right)\\ \mbox{\sc GreedyML}(\ell-1,id)\end{cases}&id\text{ mod }b^{\ell}=0\\ \text{undefined}&\text{otherwise}\end{cases}$
Figure 2: The recurrence relation for the multilevel GreedyML which is defined for each node in the accumulation tree. We denote the random subset assigned to machine $id$ by $P_{id}$.
Recurrence relation. Figure 2 shows the recurrence relation that forms the basis of the GreedyML  algorithm, defined for every node in the accumulation tree; it will be the basis for the multilevel distributed algorithm. At level 00 (leaves), the recurrence function returns the Greedy solution of the random subset of data $P_{id}$ assigned to it. At other levels (internal nodes), it returns the better among the Greedy solution computed from the union of the received solution sets of its children and its solution from its previous level. It is undefined for $(\ell,id)$ tuples that do not correspond to nodes in the tree (at higher levels). The detailed pseudocode of our algorithm is presented in Algorithm 2 in Appendix B.
4 Analysis of Our Algorithm
We prove the expected approximation ratio of GreedyML algorithm in Theorem 4.4 using three Lemmas. We restate a result from [2] that applies to the leaves of our accumulation tree and characterizes elements that do not change the solution computed by the Greedy algorithm.
Lemma 4.1 ([2]).
If we have Greedy $(V\cup\{e\})=\textsc{Greedy}(V)$, for each element $e\in B$, then $\textsc{Greedy}(V\cup B)=\textsc{Greedy}(V)$.
The next two Lemmas connect the quality of the computed solutions to the optimal solution at the internal nodes (in level one) of the accumulation tree. Lemma 4.2 provides a lower bound on the expected function value of the individual solutions of the Greedy algorithm received from the leaf nodes while Lemma 4.3 analyzes the expected function value of the Greedy execution over the accumulated partial solutions. Due to space constraints, we state the lemmas here and provide the proofs in the Appendix D.
Let $p\colon V\to[0,1]$ be a probability distribution over the elements in $V$, and $A\sim V(1/m)$ be a random subset of $V$ such that each element is independently present in $A$ with probability $1/m$. The probability $p$ is defined as follows:
$p(e)=\begin{cases}\Pr\limits_{A\sim V(1/m)}\left[e\in\textsc{Greedy}(A\cup\{e\})\right],&\text{if}\ e\in OPT;\\ 0,&\text{otherwise.}\end{cases}$
For any leaf node, the distribution $p$ defines the probability that each element of $OPT$ is in the solution of the Greedy algorithm when it is placed in the node.
Lemma 4.2.
Let $c$ be a leaf node of the accumulation tree, $S_{c}$ be the solution computed from $c$, and $V_{c}\subset V_{n}$ be the elements considered in forming $S_{c}$. If Greedy is an $\alpha$-approximate algorithm, then ${\mathbb{E}}_{V_{n}}[f(S_{c})]\geqslant\alpha\cdot\widehat{f}(1_{OPT}-p).$
Lemma 4.3.
Let $D$ be the union of all the solutions computed by the $b$ children of an internal node $(1,id)$ in the accumulation tree, and $S$ be the solution from the Greedy algorithm on the set $D$. If Greedy is an $\alpha$-approximate algorithm, then $\mathbb{E}_{V_{n}}[f(S)]\geqslant\dfrac{b*\alpha}{m}\cdot\widehat{f}(p).$
Theorem 4.4.
Let $T(m,L,b)$ be an accumulation tree, $V$ be the ground set, and $r_{V}$ be a random mapping of elements of $V$ to the leaves of the tree $T$. Let $OPT$ be an optimal solution computed from $V$ for the constrained submodular function $f$. If Greedy is an $\alpha$-approximate algorithm, then ${\mathbb{E}\left[f(\mbox{\sc GreedyML}(V))\right]\geqslant\dfrac{b\cdot\alpha}{(m+b)}f(OPT).}$
Proof 4.5.
We concentrate on a node at level 1, where after obtaining the partial solutions from the children of this node, we compute the Greedy on the union of these partial solution. Let $S_{c}$ be any of the partial solutions, $S$ be the union of these partial solutions, and $T$ be $\arg\max\{f(S),f(S_{c})\}$. From Lemma 4.2 and Lemma 4.3,
$\mathbb{E}[f(T)]\geqslant\frac{b\alpha}{m}\cdot\widehat{f}(p)\text{ and }\mathbb{E}[f(T)]\geqslant\alpha\cdot\widehat{f}(1_{OPT}-p).$
By multiplying the first inequality by $m/b$ and then adding it to the second, we get
[…]
$\mathbb{E}[f(T)]$ $\geqslant\frac{\alpha}{(m/b+1)}\cdot\widehat{f}(1_{OPT}).$
The theorem follows since the solution quality can only improve at higher levels of the tree.
5 Experimentation
Experimental Setup.
We conduct experiments to evaluate our algorithms using different accumulation tree structures and compare them with Greedy and RandGreedi to assess the quality, runtime, and memory footprints of these algorithms. All the algorithms are executed on a cluster computer with $448$ nodes, each of which is an AMD EPYC $7662$ node with $256$ GB of total memory shared by the $128$ cores. Each core operates at $2.0$ GHz frequency. To simulate a completely distributed environment on this cluster, we needed to ensure that the memory is not shared between nodes. Therefore, in what follows, a machine will denote one node with just one core assigned for computation, but having access to all $256$ GB of memory. We also found that this made the run time results more reproducible.
For our experimental evaluation, we report the runtime and quality of the algorithms being compared. For runtime, we exclude the file reading time in each machine, and for the quality, we show the objective function value of the corresponding submodular function. Since the RandGreedi and GreedyML are distributed algorithms, we also report the number of function calls in the critical path of the computational tree, which represents the parallel runtime of the algorithm. Given an accumulation tree, the number of function calls in the critical path refers to the maximum number of function calls that the algorithm makes along a path from the leaf to the root. In our implementation, this quantity can be captured by the number of function calls made by the nodes of the accumulation tree with $id=0$ since this node participates in the function calls from all levels of the tree.
Function Dataset $n=|V|$ $\sum_{u}\delta(u)$ avg. $\delta(u)$
$k$-dominating set AGATHA_2015 183,964,077 11,588,725,964 63.32
MOLIERE_2016 30,239,687 6,669,254,694 220.54
com-Friendster 65,608,366 1,806,067,135 27.52
road_usa 23,947,347 57,708,624 2.41
road_central 14,081,816 33,866,826 2.41
belgium_osm 1,441,295 3,099,940 2.14
$k$-cover webdocs 1,692,082 299,887,139 177.22
kosarak 990,002 8,018,988 8.09
retail 88,162 908,576 10.31
$k$-medoid Tiny ImageNet 100,000 1,228,800,000 12,288
Table 1: Properties of Datasets used in the experiments. $\delta(u)$ is the number of neighbors of vertex $u$ for the $k$-dominating set problem, the cardinality of the subset $u$ for the $k$-cover problem, and the size of the vector representation of the pixels of image $u$ for the $k$-medoid problem.
Datasets. In this paper, we limit our experiments to cardinality constraints using three different submodular functions described in detail in Appendix C.
Our benchmark dataset is shown in Table 1. They are grouped based on the objective function and are sorted by the $\sum_{u}\delta(u)$ values within each group (see the Table for a definition). For the $k$-dominating set, our testbed consists of the Friendster social network graph [29], a collection of road networks from DIMACS10 dataset and the Sybrandt dataset. We chose road graphs since they have relatively small average vertex degrees, leading to large vertex-dominating sets. We chose the Sybrandt collection [25][24] since it is a huge data set of machine learning graphs. For the $k$-cover objective, we use popular set cover datasets from the Frequent Itemset Mining Dataset Repository [8]. For the $k$-medoid problem, we use the Tiny ImageNet dataset [16].
MPI Implementation. GreedyML is implemented using C++11, and compiled with g++9.3.0, using the O3 optimization flag. We use the Lazy Greedy [20] variant that has the same approximation guarantee as the Greedy but is faster in practice since it potentially reduces the number of function evaluations needed to choose the next element (by using the monotone decreasing gain property of submodular functions). Our implementation of the GreedyML algorithm uses Open MPI implementation for the inter-node communication. We use the MPI_Gather and MPI_Gatherv primitives to receive all the solution sets from the children (Line 11 in Algorithm 2). We generated custom MPI_Comm communicators to enable this communication using MPI_Group primitives. Customized communicators are required since every machine has different children at each level. Additionally, we use the MPI_Barrier primitive to synchronize all the computations at each level.
5.1 Experimental Results
The experiments are executed with different accumulation trees that vary in the number of machines ($m$), the number of levels $(L)$, and the branching factors $(b)$, in order to assess their performance. We repeat each experiment six times and report the geometric mean of the results. Unless otherwise stated, a machine in our experiments represents a node in the cluster with only one core assigned for computation.Whenever memory constraints allow, we compare our results with the sequential Greedy  algorithm that achieves $(1-1/e)$ approximation guarantee.
Recall that our GreedyML algorithm generalizes the RandGreedi algorithm by allowing multiple levels in the accumulation tree, thus removing the bottleneck of a single aggregation. In the following, we verify this through a series of experiments.
In Section 5.1.1, we assess the performance of our algorithm using different accumulation tree structures. We fix the number of machines and construct the best parameters of the accumulation tree for our dataset. Additionally, the experiment also demonstrates that the number of function calls in the critical path is a good estimate of the parallel runtime. In Section 5.1.2, we show the memory benefit of our GreedyML w.r.t RandGreedi with two experiments. In Section 5.1.2, we impose a limit of 100 MB space for each node and vary $k$, the selection size. This also simulates how the new algorithm can find applications in the edge-computing context. In Section 5.1.2, we fix $k$ and vary the memory limits, necessitating different numbers of nodes to fit the data in the leaves. We observe the quality and runtime of different accumulation tree structures in these two experiments. Both these experiments are designed to show that the RandGreedi algorithm quickly runs out of memory with increasing $m$ and $k$, and by choosing an appropriate accumulation tree, our GreedyML algorithm can solve this problem with negligible drop in accuracy. For these experiments, we will choose the computational tree with the lowest depth that can be used with the memory limit and $k$ values.
In Section 5.1.3, we perform a scaling experiment by varying the number of machines and using the tallest tree by setting a branching factor of two for the accumulation tree. We specifically show that even though the RandGreedi algorithm has a low asymptotic communication cost, it can become a bottleneck when scaled to a large number of machines. We also show how our algorithm alleviates this bottleneck. Finally, in Section 5.1.4, we perform experiments for the k-medoid objective function and show that we can provide a significant speedup by using taller accumulation trees without loss in quality. The k-medoid function is extensively used in machine learning as a solution to exemplar-based clustering problems.
5.1.1 Accumulation tree parameter selection
Figure 3: Geometric means of results from GreedyML for $k$-dominating set (on road datasets) and for $k$-cover (on set cover benchmarks) using $32$ machines. The $L$ and $b$ represent different accumulation tree configurations. The bottom right plot shows the geometric means of the number of function calls in the critical path relative to the Greedy algorithm for $k=32,000$.The remaining three subfigures show the execution times for different $k$ values and accumulation trees.
In this experiment, we show results for the $k$-dominating set and $k$-coverage problem by fixing the number of machines and varying branching factors, the number of levels in the accumulation tree, and the selection set size $k$. In Figure 3, we provide summary results on the number of function evaluations in the critical path relative to the Greedy algorithm and the running times by taking a geometric mean over all nine datasets.
Three subfigures (top left, top right, and bottom left) of Figure 3 show the execution time in seconds for the GreedyML and RandGreedi algorithms, as the number of levels and the parameter $k$ are varied. When $k$ is small (top left), there is less variation in the execution time since work performed on the leaves dominates overall time. As $k$ increases (bottom left), the GreedyML algorithm becomes faster than the RandGreedi algorithm ($L=1,b=32$). Note that although Figure 3 presents the geometric mean results over all nine datasets, the runtime and the function values for the individual datasets follow the same trend. The largest and smallest reduction in runtime we observe is on the belgium_osm and kosarak datasets with a reduction of around $22\%$ and 1% across, respectively, for all $k$ values.
The bottom right plot fixes $k=32,000$ and shows the number of function calls in the critical path of the accumulation tree relative to the Greedy algorithm for different $(L,b)$ pairs. Here, the leftmost bar represents the RandGreedi algorithm. We observe that the relative number of function calls for RandGreedi is around 70% of Greedy, whereas the GreedyML (with $L=2$ and $b=8$) reduces it by 15 percent. From Table 4, the function calls complexity at a leaf node is $O(nk/m)$ and at an accumulation node, it is $O(mk^{2})$ for the RandGreedi algorithm. Hence, the accumulation node dominates the computation since it has a quadratic dependence on $k$, becoming a bottleneck for large $k$ values. This plot also shows that the number of function calls is a good indicator of the algorithm’s run time and that the cost of function evaluations dominates the overall time. The other factor affecting run time is communication costs, which are relatively small and they grow with the number of levels when $k$ is very large.
We note (not shown in the figure) that the objective function values obtained by the GreedyML algorithm are not sensitive to the choice of the number of levels and the branching factors of the accumulation tree and differ by less than $1\%$ from the values of the RandGreedi algorithm. For the webdocs $k$-coverage problem, however, Greedy quality is about $20\%$ higher than both the RandGreedi and GreedyML.
In Table 6 in Appendix 7, we report the advantages of accumulating in multiple levels over choosing to stop at level one of accumulation. We use maximum $k$-cover function for the Friendster dataset with $k=1000$ for different branching factors at the first and final accumulation levels. We observe the objective values at the highest accumulation level are not very sensitive to the tree parameters, contrary to their sensitivity to the approximation ratio derived in Theorem 4.4.
Figure 4: Results from GreedyML for the $k$-dominating set problem on the road_usa dataset on 16 nodes with varying $k$. The pair($L$,$b$) shows the number of levels and branching factors chosen for specific $k$ values. The function values are relative to the Greedy algorithm. Note that the leftmost bars in both plots represent the RandGreedi results.
5.1.2 Experiments with memory limit
Varying k.
For this experiment, we use 16 machines with a limit on the available memory of 100 MB per machine and vary $k$ from $128,000$ to $1,024,000$ for the $k$-dominating set problem on the road_usa [1] dataset. The small memory limit in this experiment can also be motivated from an edge computing context.
The left plot in Figure 4 shows the number of function calls with varying values of $k$ for the Greedy (green bars) and GreedyML algorithms (blue bars). For the GreedyML (and the RandGreedi), we are interested in the number of function calls in the critical path since it represents the parallel runtime of the algorithm. With our memory limits, only $k=128,000$ instance can be solved using the RandGreedi algorithm.
As we increase $k$, we are able to generate solutions using our GreedyML with different accumulation trees. The corresponding lowest-depth accumulation tree with the number of levels and branching factor is shown on top of the blue bars. The result shows that the number of function evaluations on the critical path in the GreedyML algorithm is smaller than the number of function evaluations in the sequential Greedy algorithm. While the number of function calls for accumulation trees with smaller $b$ values is larger than RandGreedi, we see that GreedyML can solve the problems with larger $k$ values in the same machine setup, which was not possible with RandGreedi. But it comes with a trade-off on parallel runtime. We observe that as we make the branching factor smaller, the number of function calls in the critical path increases, suggesting that it is sufficient to choose the accumulation trees with the largest branching factor (thus the lowest depth tree) whenever the memory allows it.
The right plot of Figure 4 shows the relative objective function value, i.e., the relative number of vertices covered by the dominating set compared to the Greedy algorithm, with varying $k$. The figure shows that the RandGreedi and GreedyML algorithms attain quality at most $6\%$ less than the serial Greedy algorithm. Similar trends can be observed for other datasets.
Dataset Alg. Mem. Limit $m$ $b$ $L$ Rel. Func.(%) Time (s.)
Friendster RG 4GB 8 8 1 99.959 61.994
GML 2GB 16 4 2 99.903 61.352
GML 1GB 32 2 5 99.793 79.997
MOLIERE_2016 RG 8GB 8 8 1 99.257 121.318
GML 4GB 16 4 2 99.106 108.764
GML 2GB 32 2 5 98.990 161.139
AGATHA_2015 RG 12GB 8 8 1 99.996 94.122
GML 6GB 16 4 2 99.995 99.574
GML 3GB 32 2 5 99.989 104.156
Table 2: Results for $k$-dominating set on the Friendster, road_usa and webdocs datasets. The memory size per machine is varied for the Friendster dataset. The number of machines $m$ and the accumulation tree are selected based on the size of the data and the size of the solutions to get three different machine configurations. We report the function values relative to the Greedy algorithm and the execution time in seconds. Note that the 4GB entries run with $L=1$ and correspond to the RandGreedi (RG) algorithm. We use the same three machine organizations for the road_usa and webdocs datasets to show they follow similar trends in solution quality and execution time.
Varying memory limits.
This experiment demonstrates that the memory efficiency of the GreedyML algorithm enables us to solve problems on parallel machines, whereas the RandGreedi and Greedy cannot solve them due to insufficient memory. Unlike the previous experiment (Varying $k$), where we selected the accumulation trees based on $k$, here, we fix $k$ and choose accumulation trees based on the memory available on the machines. We consider the $k$-dominating set problem and report results on the Friendster [29], AGATHA_2015[25], and MOLIERE_2016[24] dataset in Table 2. For the Friendster dataset, we choose $k$ such that the $k$-dominating set requires $512$ MB, roughly a factor of $64$ smaller than the original graph. The RandGreedi algorithm (the first row) can execute this problem only on $8$ machines, each with $4$ GB of memory, since in the accumulation step, one machine receives solutions of size $512$ MB each from $8$ machines. The GreedyML algorithm having multiple levels of accumulation can run on $16$ machines with only $2$ GB memory, using $L=2$ and $b=4$. Furthermore, it can also run on $32$ machines with only $1$ GB memory, using $L=5$ and $b=2$. We repeat the same experiment for the other two datasets with these three machine configurations with corresponding memory restrictions.
We show relative quality and running time for the three datasets from these configurations in Table 2. Our results show that function values computed by the GreedyML algorithm (the $2$ and $1$ GB results) are insensitive to the number of levels in the tree. As expected, increasing the number of levels in the accumulation tree increases the execution times due to the communication and synchronization costs involved. However, aggregating in multiple levels enables us to solve large problems by overcoming memory constraints. So, in this scenario, it is sufficient to select the number of machines depending on the size of the dataset and then select the branching factor such that the accumulation step does not exceed the memory limits. We also notice that the RandGreedi algorithm has an inherently serial accumulation step, and the GreedyML algorithm provides a mechanism to parallelize it.
Figure 5: Strong scaling results of the RandGreedi and GreedyML algorithms for $k=50$ on Friendster dataset for the $k$-dominating set problem. We set $b=2$ for the GreedyML algorithm.
5.1.3 Scaling results
Next, we show how the GreedyML algorithm alleviates the scaling bottlenecks of the RandGreedi algorithm using the $k$-dominating set problem on the Friendster dataset. We set the branching factor $b=2$ for the GreedyML algorithm since this has the highest number of levels and, thus, the lowest approximation guarantee. We compare communication and computation times against RandGreedi algorithm from 8 to 128 machines with $k=50$.
In Figure 5, we plot the total execution time by stacking communication and computation times for the two algorithms. For RandGreedi, the communication time scales poorly since it increases linearly with the number of machines (See Table 4). But, for GreedyML algorithm (with a constant branching factor, $b=2,L=\log_{2}m$), the communication cost is $O(k\log m)$, which grows logarithmically in the number of machines. Figure 5 shows that the total communication times of the GreedyML algorithm are consistently around $0.25$ seconds, whereas the RandGreedi increases from $0.05$ second to $2$ seconds. We observe that computation times for both RandGreedi and GreedyML changes similarly with $m$, indicating that the majority of the computation work is performed at the leaf nodes. For computation time, we observe a slightly worse scaling of RandGreedi compared to GreedyML, again because the central node becomes a computational bottleneck as $m$ increases. Similar to other experiments, we observe (not shown in the plot) an almost identical quality in the solutions, where the GreedyML solution has a quality reduced by less than $1\%$ from that of the RandGreedi algorithm.
5.1.4 The k-medoid problem
$L$ $b$ Local Obj. Added Images
Rel. Func. Val. (%) Speedup Rel. Func. Val. (%) Speedup
5 2 92.22 2.00 93.69 2.01
3 4 92.21 1.96 92.70 1.94
2 8 92.73 1.95 92.77 1.93
2 16 92.22 1.49 93.34 1.44
Table 3: Results from GreedyML for the $k$-medoid function on the Tiny ImageNet data set using different accumulation trees. The table shows the relative function values and speedup compared to the RandGreedi algorithm using two different schemes for computing the local objective functions on 32 nodes. Higher values are better for both schemes. Recall that $L$ and $b$ are the number of levels and branching factor, respectively.
Our final experiment considers the $k$-medoid function that solves the exemplar-based clustering problem. Our dataset consists of the Tiny ImageNet dataset [16] containing 100K images ($64\times 64$ pixels) with $200$ different classes with $500$ images from each class. We convert and normalize the image into a vector and use Euclidean distance to measure dissimilarity. We define an auxiliary vector $e_{0}$ as a pixel vector of all zeros. Note that, unlike the other two functions, the $k$-medoid function requires access to the full dataset to compute the functional value. Since the dataset is distributed, this poses an issue in the experiment. To overcome this, following [21, 2], we calculate the objective function value using only the images available locally on each machine. This means the ground set for each machine is just the images present in that machine. Additionally, they [21, 2] have also added subsets of randomly chosen images to the central machine to provide practical quality improvement. We have followed these techniques (local only and local with additional images) in the experiments for our multilevel GreedyML algorithm.
In our experiments, we set $k$ to $200$ images, fix the number of machines ($m=32$), and vary the accumulation trees by choosing different $L$ and $b$. For the variant with additional images, we add $1,000$ random images from the original dataset to each accumulation step.
In Table 3, we show the relative objective function values and speedup for different accumulation trees relative to the RandGreedi algorithm. We observe that the objective function values for GreedyML algorithm are almost similar to RandGreedi. Our results show that the GreedyML algorithm becomes gradually faster as we increase the number of levels with runtime improvement ranging from 1.452.01×1.45-2.01$\times$1.45 - 2.01 ×. This is because the $k$-medoid function is compute-intensive, where computation cost increases quadratically with the number of images (Table 4). With $k=200$ and $m=32$, the RandGreedi algorithm has $km=6,400$ images at the root node but only $n/m=313$ images at the leaves; thus the computation at the root node dominates in cost. On the other hand, as we decrease the branching factor (from $b=16$ to $2$), the number of images ($kb$) in the interior nodes decreases from $3,200$ to $400$ for the GreedyML algorithm. This gradual decrease in compute time is reflected in the total time and in the observed speedup.
Finally, in Fig. 7 (Appendix E), we show $16$ out of the $200$ images determined to be cluster centers by the GreedyML and RandGreedi algorithms. We can conclude that the submodular $k$-medoid function can generate a diverse set of exemplar images for this clustering problem.
6 Conclusion and Future work
[…]
In the future, we plan to conduct experiments for other hereditary constraints, such as matroid and $p$-system constraints. Another direction is to apply GreedyML to closely related non-monotone and weakly submodular functions. Our experiments suggest that GreedyML delivers higher quality solutions than the expected approximation guarantees. One area of future work could involve investigating whether the approximation ratio can be further improved.
Appendix A Accumulation Trees
{forest}
for tree=circle,draw, thick, minimum size=4.1em, s sep = 1em, inner sep=1pt, l=6em [${3,0}$ [${2,0}$ [${1,0}$ [${0,0}$] [${0,1}$] ][${1,2}$ [${0,2}$] [${0,3}$] ] ][${2,4}$ [${1,4}$ [${0,4}$] [${0,5}$] ][${1,6}$ [${0,6}$] [${0,7}$] ] ] ]
{forest}
for tree=circle,draw, thick, minimum size=4.1em, s sep = 1em, inner sep=1pt, l=6em [${2,0}$ [${1,0}$ [${0,0}$] [${0,1}$] [${0,2}$] ][${1,3}$ [${0,3}$] [${0,4}$] [${0,5}$] ][${1,6}$ [${0,6}$] [${0,7}$] ] ]
{forest}
for tree=circle,draw, thick, minimum size=4.1em, s sep = 1em, inner sep=1pt, l=6em [${2,0}$ [${1,0}$ [${0,0}$] [${0,1}$] [${0,2}$] [${0,3}$] ][${1,4}$ [${0,4}$] [${0,5}$] [${0,6}$] [${0,7}$] ] ]
{forest}
for tree=circle,draw, thick, minimum size=4.1em, s sep = 1em, inner sep=1pt, l=9em [${1,0}$, [${0,0}$] [${0,1}$] [${0,2}$] [${0,3}$] [${0,4}$] [${0,5}$] [${0,6}$] [${0,7}$] ]
Figure 6: Accumulation tree with 8 machines and branching factors 2 (top-left), 3 (top-right), 4 (bottom-left), and 8 (bottom-right). The labels inside a node represent the identification of the node.
Appendix B Omitted Pseudocodes
B.1 Pseudocode of  RandGreedi
Algorithm 1 RandGreedi framework for maximizing constrained submodular function
1:
2:procedure RandGreedi($V$: Dataset, $m$: number of machines)
3:    $S\leftarrow\emptyset$
4:    Let $\{P_{0},P_{1},\ldots,P_{m-1}\}$ be an uniform random partition of $V$.
5:    Run $\textsc{Greedy}(P_{i})$ on each machine $i\in[0,m-1]$ to compute the solution $S_{i}$
6:    Place $S=\bigcup_{i}S_{i}$ on machine 0
7:    Run $\textsc{Greedy}(S)$ to compute the solution $T$ on machine 0
8:    return $\arg\max\left\{f(T),f(S_{1}),f(S_{2}),\dotsc,f(S_{m-1})\right\}$
9:end procedure
B.2 Pseudocode of  GreedyML
Algorithm 2 describes our multilevel distributed algorithm using two procedures. The first procedure GreedyML is a wrapper function that sets up the environment to run the distributed algorithm. The second function $\mbox{\sc GreedyML}^{\prime}$ is the iterative implementation of the recurrence relation that runs on each machine. The wrapper function partitions the data into $m$ subsets and assigns them to the machines (Line 2). Then each machine runs the $\mbox{\sc GreedyML}^{\prime}$ function on the subset assigned to it (Line 5, Line 7). The wrapper function uses and returns the solution from machine 00 (Line 8) as it is the root of the accumulation tree.
The $\mbox{\sc GreedyML}^{\prime}$ procedure is an iterative implementation of the recurrence relation 2 that runs on every machine. Each machine checks whether it needs to be active at a particular level (Line 5) and decides whether it needs to receive from (Line 11) or send to other machines (Line 6). The function returns the solution from the last level of the machine.
Algorithm 2 Our Randomized Multi-level GreedyMLAlgorithm
1:procedure GreedyML($V$: Dataset, $b$: branching factor, $m$: number of machines, $r$: random tape)
2:    Let $\{P_{0},P_{1},\ldots P_{m-1}\}$ be uniform random partition of $V$ using $r$.
3:    for $i=1\ldots m-1\text{ in parallel}$ $\triangleright$ Run GreedyML’ on all machines except 00
4:         $\ell=level(i,b)$triangleright $level(i,b)=\max\limits_{l}\{l:id\mod b^{l}$ is 0}0\}0 }
5:         Run $\mbox{\sc GreedyML}^{\prime}(V_{i},\ell,b,i)$ to obtain $S_{i}$ on machine $i$
6:    end for
7:    Run $\mbox{\sc GreedyML}^{\prime}(V_{0},\lceil\log_{b}m\rceil,b,0)$ to obtain $S_{0}$ on machine 0
8:    return $S_{0}$
9:end procedure
1:procedure $\mbox{\sc GreedyML}^{\prime}$($P$: Partial Data-set, $\ell$: levels; $b$: branching factor, $id$: machine ID)
2:    $S=\textsc{Greedy}(P)$
3:    $S_{prev}=S$
4:    for $i=1\dots\ell$ do
5:         if $id\neq parent(id,i)$  then
6:             Send $S_{prev}$ to $parent(id,i)$triangleright $parent(id,i)=b^{i}\cdot\lfloor id/b^{i}\rfloor$
7:             break
8:         end if
9:         $D=S_{prev}$triangleright Prepare $D$ for current iteration
10:         for $j=1\dots b-1$ do
11:             Receive $D_{j}$ from $child(id,i,j)$triangleright $child(id,i,j)=id+j\cdot b^{i-1}$
12:             $D=D\cup D_{j}$
13:         end for
14:         Run $\textsc{Greedy}(D)$ to obtain $S$
15:         $S_{prev}=\arg\max\{f(S),f(S_{prev})\}$
16:    end for
17:    return $S_{prev}$
18:end procedure
Appendix C Submodular Functions and Complexity
[…]
In our experiments, we have considered the following three submodular functions.
$k$-cover
Given a ground set $B$, a collection of subsets $V\subseteq 2^{B}$, and an integer $k$, the goal is to select a set $S\subseteq V$ containing $k$ of these subsets to maximize $f(S)=|\bigcup_{S_{i}\in S}S_{i}|$.
$k$-dominating set
The $k$-dominating set problem is a special case of the $k$-cover problem defined on graphs with the ground set $V$ as the set of vertices. We say a vertex $u\in V$ dominates all its adjacent vertices (denoted by $\delta(u)$). Our goal is to select a set $S$ of $k$ vertices to dominate as many vertices as possible, i.e., $f(S)=|\bigcup_{u\in S}\delta(u)|$. The marginal gain of any vertex is the number of vertices in its neighborhood that are not yet dominated. Therefore, the problem shows diminishing marginal gains and is submodular.
$k$-medoid problem
The $k$-medoid problem [10] is used to compute exemplar-based clustering, which asks for a set of exemplars (cluster centers) representatives of a large dataset. Given a collection of elements in a ground set $V$, and a dissimilarity measure $d(u,v)$, we define a loss function (denoted by $L$) as the average pairwise dissimilarity between the exemplars ($S$) and the elements of the data set, i.e., $L(S)=\nicefrac{{1}}{{|V|}}\sum_{u\in V}\min_{v\in S}d(u,v)$. Following [21], we turn this loss minimization to a submodular maximization problem by setting $f(S)=L(\{e_{0}\}-L(S\cup\{e_{0}\}$, where $e_{0}$ is an auxiliary element specific to the dataset. The goal is to select a subset $S\subseteq V$ of size $k$ that maximizes $f(S)$.
Next, we analyze the computational and communication complexity of our GreedyML algorithm using the bulk synchronous parallel (BSP) model of parallel computation [28]. We denote the number of elements in the ground set by $n=|V|,$the solution size by $k$, the number of machines by $m$, and the number of levels in the accumulation tree by $L$.
Computational Complexity. The number of objective function calls by the sequential Greedy algorithm is $O(nk)$, since $k$ elements are selected to be in the solution, and we may need to compute $O(n)$ marginal gains for each of them. Each machine in RandGreedi algorithm makes $O(k(n/m+mk))$ function calls, where the second term comes from the accumulation step. Each machine of the GreedyML algorithm with branching factor $b$ makes $O(k(n/m+Lbk))$ calls. Recall that $L=\lceil\log_{b}{m}\rceil$.
We note that the time complexity of a function call depends on the specific function being computed. For example, in the $k$-coverage and the k-dominating set problems, computing a function costs $O(\delta)$, where $\delta$ is the size of the largest itemset for $k$-coverage, and the maximum degree of a vertex for the vertex dominating set. In both cases, the runtime complexity is $O(\delta k(n/m+mk))$ for the RandGreedi, and $O(\delta k(n/m+Lbk))$ for the GreedyML algorithm. The $k$-medoid problem computes a local objective function value and has a complexity of $O(n^{\prime}\delta)$ where $\delta$ is the number of features, and $n^{\prime}$ is the number of elements present in the machine. For the leaves of the accumulation tree, $n^{\prime}=n/m$, and for interior nodes, $n^{\prime}=bk$. Therefore its complexity is $O(k\delta((n/m)^{2}+(mk)^{2}))$ for the RandGreedi, and $O(k\delta((n/m)^{2}+L(bk)^{2}))$ for the GreedyML algorithm.
Communication Complexity. Each edge in the accumulation tree represents communication from a machine at a lower level to one at a higher level and contains four messages. They are the indices of the selected elements of size $k$, the size of the data associated with each selection (proportional to the size of each adjacency list ($\leqslant\delta$), the total size of the data elements, and the data associated with each selection. Therefore the total volume of communication is $O(k\delta)$ per child. Since at each level, a parent node receives messages from $b$ children, the communication complexity is $O(k\delta Lb)$ for each parent. Therefore the communication complexity for the RandGreedi algorithm is $O(k\delta m)$ and for the GreedyML algorithm is $O(k\delta L\left\lceil{m^{1/L}}\right\rceil).$ We summarize these results in Table 4.
Algorithms Metric Greedy RandGreedi GreedyML
All Elements per leaf node $n$ $n/m$ $n/m$
Calls per leaf node $nk$ $nk/m$ $nk/m$
Elements per interior node 00 $km$ $k\left\lceil{m^{1/L}}\right\rceil$
Calls per interior node 00 $k^{2}m$ $k^{2}\left\lceil{m^{1/L}}\right\rceil$
Total Function Calls $kn$ $k(n/m+km)$ $k(n/m+Lk\left\lceil{m^{1/L}}\right\rceil)$
$k$-cover / k-dominating set $\delta$:subset size/number of neighbours
Cost Per call $\delta$ $\delta$ $\delta$
Computational complexity $\delta kn$ $\delta k(n/m+km)$ $\delta k(n/m+Lk\left\lceil{m^{1/L}}\right\rceil)$
Communication cost 0 $\delta km$ $\delta kL\left\lceil{m^{1/L}}\right\rceil$
$k$-medoid $\delta$: number of features
Cost Per call in Leaf node $\delta n$ $\delta n/m$ $\delta n/m$
Cost Per call in interior node 0 $\delta km$ $\delta k\left\lceil{m^{1/L}}\right\rceil$
Computational complexity $\delta kn^{2}$ $\delta k((n/m)^{2}+(km)^{2})$ $\delta k((n/m)^{2}+L(k\left\lceil{m^{1/L}}\right\rceil)^{2})$
Communication cost 0 $\delta km$ $\delta kL\left\lceil{m^{1/L}}\right\rceil$
Table 4: Complexity Results of the submodular functions for different algorithms. The number of elements in the ground set is $n$, the selection size is $k$, the number of machines is $m$, and the number of levels in the accumulation tree is $L$.
Parameter Description Parameter Description
$\alpha$ Approximation ratio of the Greedy algorithm $W$ Complete universe for the input dataset
$b$ Branching factor of the accumulation tree $w$ Size of the universe $W$
$m$ Number of leaves in the accumulation tree $V$ Input Dataset
Numbers of machines used for computation. $n$ Size of the input dataset $V$
$L$ Number of levels of the accumulation tree $V_{c}$ Dataset corresponding to any node $c$ of the tree
$\ell$ level identifier for a node $S$ Solution Set
$id$ Machine identifier for a node. $k$ Size of solution
$f$ Submodular function $OPT$ The optimal solution
$\widehat{f}$ Lovász extension of function $f$ $p(e)$ Probability that e $\in$ OPT is selected
$P_{id}$ Part of the dataset assigned to machine $id$ by the Greedy algorithm when sampled from V.
Table 5: Notations and parameters used in the paper.
Appendix D Omitted Proofs
D.1 Proof of Lemma 4.2
Proof D.1.
We first construct a subset of $OPT$ that contains all the elements that do not appear in $S_{c}$ when added to some leaf node in the subtree rooted at child $c$. Let $O_{c}$ be the rejected set that can be added to $V_{c}$ without changing $S_{c}$; i.e., $O_{c}=\{e\in OPT:e\notin\textsc{Greedy}(V_{c}\cup\{e\})\}.$ Therefore, $\Pr[e\in O_{c}]=1-\Pr[e\notin O_{c}]=1-p(e).$
From Lemma 4.1, we know that $\textsc{Greedy}(V_{c}\cup O_{c})=\textsc{Greedy}(V_{c})$ Since the rejected set $O_{c}\subseteq OPT$ and the constraints are hereditary, $O_{c}\in\mathcal{C}$ (i.e $O_{c}$ is a feasible solution of child node c). Then from the condition of Lemma 4.2, we have
$f(S_{c})$ $\geqslant\alpha\cdot f(O_{c})$
$\mathbb{E}[f(S_{c})]$ $\geqslant\alpha\cdot\mathbb{E}[f(O_{c})]=\alpha\cdot\widehat{f}(\mathbb{E}_{V_{n}}[1_{O_{c}}])=\alpha\cdot\widehat{f}(1_{OPT_{\ell,id}}-p_{\ell,id}).$
D.2 Proof of Lemma 4.3
Proof D.2.
We first show a preliminary result on the union set $D$. Consider an element $e\in D\cap OPT$ present in some solution $S_{c}$ from a child $c$. Then,
$\Pr[e\in S_{c}|e\in V_{c}]=\Pr[e\in\textsc{Greedy}(V_{c})|e\in V_{c}].$
Since the distribution of $V_{c}\sim V(1/m)$ conditioned on $e\in V_{c}$ is identical to the distribution of $B\cup\{e\}$, where $B\sim V(1/m)$, we have,
$\Pr[e\in S_{c}|e\in V_{c}]=\Pr_{B\sim V(1/m)}[e\in\textsc{Greedy}(B\cup\{e\})]=p(e).$
Since this result holds for every child $c$, and each subset $V_{c}$ is disjoint from the corresponding subsets mapped to the other children, we have
$\Pr(e\in D\cap OPT)$ $=\sum_{i}\Pr[e\in S_{c_{i}}\cap OPT|e\in V_{c_{i}}]\Pr[e\in V_{c_{i}}].$ $=\sum_{i}p*1/m=bp/m.$
Now, we are ready to prove the Lemma. The subset $D\cap OPT_{\ell,id}\in\mathcal{C}$, since it is a subset of $OPT_{\ell,id}$ and the constraints are hereditary. Further, since the Greedy algorithm is $\alpha$-approximate, we have
$f(S)$ $\geqslant\alpha\cdot f(D\cap OPT_{\ell,id})$
$\mathbb{E}_{V_{n}}[f(S)]$ $\geqslant\mathbb{E}_{V_{n}}[\alpha\cdot f(D\cap OPT)]$
[…]
Appendix E Omitted Results
Figure 7: Results from GreedyML for the $k$-medoid problem on the Tiny ImageNet dataset on 32 nodes with $k=200$ with no images added at each accumulation step. The subfigure on the left shows the first 16 image results for one of the runs for the GreedyML algorithm with branching factor $b=2$, and the subfigure on the right shows the top 16 image results for one of the runs for the RandGreedi algorithm.
Branching Factor First accumulation Final accumulation
4 0.74111 0.99994
8 0.82971 1.00005
16 0.91965 0.99994
32 1.00003 1.00003
Table 6: Objective function values relative to the Greedy algorithm at the first and final accumulation steps for Friendster with selection size $k=1000$ and $m=32$.